OpenStudy (anonymous):

In triangle STU, s = 9 cm, t = 15 cm, and m angle U = 37 degrees. Find u.

OpenStudy (raden):

 use the cosine rule : u^2 = s^2 + t^2 - 2 * s * t * cosU u^2 = 9^2 + 15^2 - 2 * 9 * 15 cos37 solve for u (simplify and use the calculator)